OpenStudy (anonymous):

find the LCM of 62,210 and 75

OpenStudy (anonymous):

32550

OpenStudy (anonymous):

pls how did u calculate??

OpenStudy (anonymous):

Do you know how to get prime factors?

OpenStudy (anonymous):

Kawsar i think i can,using the table formula

OpenStudy (anonymous):

It's not 32250, that's not a multiple of 62210. To get LCM, factor each number into its primes: 62210 = 2 * 5 * 62210 75 = 3 * 5 * 5 Then take all unique factors, and for factors that repeat, only consider one of them, (which is why I typed it out this way, so you can see 5 repeats) , i.e 2 * 3 * 5 * 5 * 62210 = 933150 LCM = 933150

OpenStudy (anonymous):

Oh, I thought the question was 62, 210 and 75, as in 3 numbers.

OpenStudy (anonymous):

Oh I didn't think of that...apologies, it very well could be. Which is it, Cuira?

OpenStudy (anonymous):

I just assumed it was 3 numbers as 62210 seemed a very high number. It could be any of the two. lol!

OpenStudy (anonymous):

oh sorry for the confusion but it is three seperate numbers

OpenStudy (anonymous):

i tried and also got 32550

OpenStudy (anonymous):

pls help me with theses ones too: lcm of 510, 210,735, and 365